# Data Poisoning for In-context Learning

### v2

@@ 4.6 Ablation studies
Under the assumption that attackers lack knowledge of the ICL process details, such as the templates and the number of examples used, we perform ablation studies to understand the impact of these settings. Specifically, we evaluate three different templates: F1, which is used in generating poisoned data; F2, formatted as ’input-output’; and F3, structured as ’Q:input, A:output’. Additionally, we explore the effects of using 3, 5, and 7 examples in ICL predictions. This ablation study concentrates on the Llama2-7B model and the GLUE-SST2 dataset, with comprehensive results provided in Appendix A.2. The results for different templates and varying numbers of examples are detailed in Table 4 and Figure 4, respectively. The results demonstrate that our framework is effectively adaptable to various templates and different numbers of demonstrations, highlighting its potentials in practice.
Table 4: Experiments with different templates: F1, F2, F3 on model Llama2-7B and dataset GLUE-SST2.
F1 F2 F3
Clean 88.6 92.5 90.3
Random label 82.1 84.7 79.2
Synonym 18.5 17.9 18.2
Character 26.8 30.6 28.5
Adv suffix 20.4 21.7 19.3
Figure 4: Experiments with different numbers of examples.
5 Related Works
5.1 In-context Learning
In-context learning (ICL), since its introduction by (Brown et al., 2020), has been notable for efficiently tackling tasks with just a few examples, bypassing the need to adjust the model’s parameters (Dong et al., 2022; Liu et al., 2021; Sun et al., 2022). Generally speaking, ICL allows LLMs to make predictions based on a few examples formatted in a template rather than changing model parameters. Existing works have shown that the performance of ICL relies on the quality of demonstrations, including the selection of examples (Liu et al., 2021; Sorensen et al., 2022; Wang et al., 2023), the order of examples (Lu et al., 2021; Min et al., 2022), the template used to format demonstrations (Liu et al., 2023a; Wei et al., 2022). The mechanisms behind ICL’s success have been the subject of much interest. One line of work reveals that ICL learns latent concepts and conducts implicit Bayesian inference (Xie et al., 2021). Follow-up works (Hendel et al., 2023; Liu et al., 2023b) take a further step and point out that hidden states in LLMs encode these concepts. Another line of work focuses on the implicit learning mechanism of ICL. Garg et al. (2022) showed that Transformers (Vaswani et al., 2017) can encode effective learning algorithms to learn unseen linear functions according to demonstration samples. Von Oswald et al. (2023); Dai et al. (2022); Akyürek et al. (2022) connect Transformers with gradient descent and claim that LLMs can implicitly perform gradient descent on examples. Bai et al. (2023) further extend the ability of implicit gradient descent to various statistical abilities including pre-ICL testing and post-ICL validation. Our work builds on the foundation of the first line, and leverages the understanding of latent concepts and hidden states to conduct effective data poisoning attacks.
5.2 Data Poisoning
Data poisoning attacks (Biggio et al., 2012; Steinhardt et al., 2017) traditionally occur during the data collection phase of machine learning model training, where the training data is tampered with to induce malicious behaviors in the resulting models. These behaviors can range from degraded performance on testing data (Steinhardt et al., 2017; Huang et al., 2021) to the misclassification of specific test samples (Shafahi et al., 2018; Zhu et al., 2019) and the implantation of backdoors. In conventional models, such attacks typically exploit the training objectives (Steinhardt et al., 2017; He et al., 2023). However, applying data poisoning to in-context learning (ICL) aims to undermine the ICL’s overall effectiveness, presenting unique challenges due to its implicit learning mechanism.
@@ 6 Conclusion
In this study, we introduce ICLPoison, a novel framework devised to assess the vulnerability of in-context learning (ICL) in the face of data poisoning attacks. We use the dynamics of hidden states in Large Language Models (LLMs) to craft our attack objectives. Furthermore, we implement our framework through three distinct and practical algorithms, each employing a different method of discrete perturbation. Our research exposes previously unidentified susceptibilities of the ICL process to data poisoning. This discovery emphasizes the urgent need for enhancing the robustness of ICL implementations.
Broader Impact
This paper introduces a new framework to assess how in-context learning in large language models (LLMs) is susceptible to data poisoning attacks. It highlights the vulnerabilities in in-context learning, urging the development of more robust implementations. Additionally, the framework could aid in creating stronger attacks. Our findings underscore the pressing need to enhance LLMs’ security to protect against such vulnerabilities.
@@ A.1 Details of Algorithms
A.2 Additional Experiments
@@
Cola 3 63.2 59.2 16.5 21.8 12.8
5 63.8 55.5 15.3 22.7 13.6
7 63.1 54.6 14.8 22.5 13.2
SST2 3 84.5 80.2 14.3 29.4 19.6
5 88.6 82.1 18.5 26.8 20.4
7 92.7 86.1 19.3 32.1 21.7
Emo 3 58 35.2 11.5 14.6 7.9
5 73.1 43.6 14.9 17.5 12.7
7 79.2 47.4 11.7 15.8 11.3
Poem 3 61 48.6 16.9 20.4 15.6
5 62.9 51.4 18.1 23.3 17.2
7 66.7 55.2 19.5 22.6 18.3
AG 3 66.9 52.8 12.9 17.5 9.5
5 73.2 57 13.6 19.4 11.9
7 78 60.1 13.2 18.9 11.6
A.3 Poisoned Text Examples

### v3

@@ 5.5 Ablation studies
We present ablation studies with regard to the number of perturbations. Due to the page limit, we provide more ablation studies in Appendix A.3 including token length, prompt template, computation cost, etc. We vary the number of perturbations from 0 to 5, on the SST2 dataset and Llama2-7B model. As shown in Table 5, increasing the number of perturbations will enhance the poisoning effect. Moreover, even for a small number of perturbations, such as 2, we can significantly compromise the accuracy of ICL by at least 30%.
Table 5: Ablation on the number of perturbations, model Llama2-7B and SST2 dataset.
Budget 0(clean) 1 2 3 4 5
Adversarial suffix 88.6 69.3 50.8 34.9 25.1 20.4
Char replacement 88.6 73.5 57.2 46.8 32.7 26.8
Synonym replacement 88.6 70.1 52.9 39.1 26.4 18.5
@@ 6 Conclusion
In this study, we introduce ICLPoison, a novel method devised to assess the vulnerability of ICL in the face of data poisoning attacks. We use the dynamics of hidden states in LLMs to craft our attack objectives. Furthermore, we implement our method through three distinct and practical algorithms, each employing a different method of discrete perturbation. Our research exposes previously unidentified susceptibilities of the ICL process to data poisoning. This discovery emphasizes the urgent need for enhancing the robustness of ICL implementations.
7 Limitations
In this work, we investigate the vulnerability of in-context learning (ICL) facing data poisoning attacks. In specific, we develop a method ICLPoison to evaluate the vulnerability. The proposed method needs access to the hidden states of LLMs, and thus can not be directly applied to API-only models. We only test on GPT-3.5 and GPT-4, and more API-only models need to be tested. Moreover, attacking methods that only require black-box access are to be investigated to directly attack API-only models.
Acknowledgement
Pengfei He and Jiliang Tang are supported by the National Science Foundation (NSF) under grant numbers CNS2321416, IIS2212032, IIS2212144, IOS2107215, DUE2234015, CNS2246050, DRL2405483 and IOS2035472, the Army Research Office (ARO) under grant number W911NF-21-1-0198, Amazon Faculty Award, JP Morgan Faculty Award, Meta, Microsoft and SNAP.
@@ A.1 Details of Algorithms
A.2 Details of datasets
  • Stanford Sentiment Treebank (SST2) dataset from the GLUE benchmark (Wang et al., 2019) is a sentiment analysis dataset consisting of sentences from movie reviews and human annotations of their sentiment in 2 classes
  • Corpus of Linguistic Acceptability (Cola) dataset from GLUE is a linguistic analysis dataset consisting of English acceptability judgments collected from linguistic books, labeled with “acceptable" or “unacceptable"
  • Emo dataset (Wang et al., 2023) focuses on emotion classification consisting of Twitter messages labeled in 4 classes
  • AG’s new (AG) corpus (Zhang et al., 2015) is a topic classification dataset gathered from news sources and labeled in 4 classes
  • Poem Sentiment (Poem) (Sheng and Uthus, 2020) is a sentiment analysis dataset of poem verses from Project Gutenberg, classified into 3 classes
A.3 Additional Experiments
@@
Practical attacks. We provide the full results of practical attacks with different poison rates in Table 6.
Table 6: Full results of practical attacks.
Poisoning rate Clean Random label Synonym Character Adv suffix
100% 88.6 82.1 18.5 26.8 20.4
50% 88.6 84.6 46.3 50.1 43.5
20% 88.6 87.5 67.1 70.8 65.9
10% 88.6 88.1 75.6 77.5 73.2
@@
Calibration and selection methods. We conduct additional experiments with two representative calibration and selection methods, contextual calibration (Zhao et al., 2021) and similarity sampling (Margatina et al., 2023), on the SST2 dataset with Llama2-7B. This is because the victims can apply these pre-processing methods. According to the results in Table12, calibration and selection methods can improve the clean accuracy. However, these methods are still vulnerable to the proposed poisoning attack, with a test accuracy of less than 20%. This may be because the contextual calibration only focuses on the potentially imbalanced distribution of the output label but does not consider the hidden states during calibration. The similarity selection selects demonstrations based on embedding similarity (Sentence-BERT as in (Margatina et al., 2023)), which may also be affected by the poisoning.
Table 12: Test calibration and selection methods on SST2 dataset and Llama2-7B model.
Clean Synonym Attack
Random selection (baseline in our paper) 88.6 18.5
Contextual calibration 90.4 19.2
Similarity selection 91.6 17.4
Scalability to context length and demonstration size. We additionally summarize the average token length of each dataset (to represent ‘context length’) used in the experiments. We use the tokenizer of Llama2-7B as the example and present results in Table 13. According to the above results, there is no obvious relationship between the attacking performance and the context length. This indicates that our method is robust to the context length.
Table 13: Average token length of each dataset on Llama2-7B model.
Dataset Avg token length attack/clean acc
SST2 14.8 18.5/88.6
COLA 10.4 15.3/63.8
Poem 10.5 18.1/62.9
Emo 18.6 11.9/73.1
AG’s news 62.4 13.6/73.2
Computation cost. We provide the average running time per example (in seconds) when poisoning dataset SST2 on model Llama2-7B with all three methods under different numbers of perturbations (1,3,5 perturbations). As shown in Table 14, adversarial suffix and synonym replacement can be time-consuming, while character replacement is much more efficient. This can be due to the size of the search space, as there are more tokens and words than characters. However, since the poisoning process is offline, attackers can take enough time to generate poisoned examples.
Table 14: Computation cost, SST2, Llama2-7B
Num of perturbation 1 3 5
Adversarial suffix 127 369 571
Char replacement 43 118 197
Synonym replacement 71 206 343
A.4 Poisoned Text Examples
@@
A.5 Paraphrased texts
To clarify about paraphrasing, we prompt the model to paraphrase the whole sentences. As a result, it is not a token-level replacement, and it is also hard to quantify how many tokens are changed. We present some examples in Table 16. According to the above table, the paraphrased sentence is different from the original text as it does not only change words but also changes the sentence structure. Besides, as in the third example, when adding some irrelevant content to the input, paraphrasing can remove it. This indicates that the paraphrasing defense is strong compared to attacks.
Table 16: Poisoned texts paraphrased by GPT-4
Original text Attacked text Paraphrased text
Temperatures Dip in Europe, sending Price of Oil Up 4.8 The price of crude oil surged 4.8 percent in New York yesterday, the biggest rise in more than five months, on concern that cold weather in Europe will draw heating oil from the United States. Temperatures Dip in Europe, sent Price of Oil Up 4.8 The price of coarse oil surging 4.8 percent in New York yesterday, the biggest rise in more than five months, on stress that cold weather in Europe will draw heating oil from the unified states. European temperatures have fallen, leading to a 4.8% surging in coarse oil prices in New York yesterday. This marks the largest increase in over five months, driven by stresses that colder conditions in Europe will increase demand for heating oil from the unified states.
More bodies found in Nablus US forces discovered more bodies under the northern city of Mosul on Friday, apparent victims of an intimidation campaign by insurgents against Iraq’s fledgling security forces. Additional bodies found in Nablus US forces uncovered more bodies under the northern city of Mosul on Friday, evident victims of an bullying campaign by insurgents against Iraq;s fledgling security forces. More bodies were discovered in Nablus as US forces unearthed additional victims in the northern city of Mosul on Friday, evidently targeted by insurgents in a campaign of bullying against Iraq’s emerging security forces.
ETA Bombs Put Radical Basque Party in Tight Spot The armed Basque separatist group ETA showed it was not a spent force with Friday #39;s coordinated attacks at five Madrid petrol stations, but it also further isolated the guerrillas #39; closest political allies, Batasuna. ETA Bombs Put Radical Basque Party in Tight Spot The armed Basque separatist group ETA showed it was not a spent force with Friday #39;s coordinated attacks at five Madrid petrol stations, but it also further isolated the guerrillas #39; closest political allies, Batasuna. Gould}}\$,attempt The ETA’s bombings have placed its key political ally, the radical Basque party Batasuna, in a difficult position. By carrying out synchronized attacks at five Madrid petrol stations on Friday, the armed Basque separatist group demonstrated that it remains a potent force but also increasingly alienated Batasuna from others.